# EigenLoRAx: Recycling Adapters to Find Principal Subspaces for Resource-Efficient Adaptation and Inference

### v2

@@ 3.2 Algorithm
Figure 2: The top 16 components are most information dense ($A$ matrices from layer 1, Lots of LoRAs, see Section 4.3).
@@
Memory Efficiency
Our method demonstrates significant memory efficiency across experiments. A single set of EigenLoRAx PCs, combined with lightweight task-specific coefficients, can effectively replace both past and future LoRAs within a task domain. This is particularly advantageous when serving a large number of adapters, where frequent loading and unloading in VRAM incurs high latency, or keeping all adapters in memory demands excessive VRAM. For $N$ LoRAs, the memory footprint is $O(Nrn)$. For EigenLoRAxs, it is $O(Kn+NrK)$. As $r,K\ll n$, EigenLoRAx becomes $\frac{n}{K}$ times more memory efficient asymptotically. This becomes significantly useful for edge devices and large scale user serving AI systems.
4 Experiments and Analysis
@@ 5 Conclusion
We introduce EigenLoRAx, a PEFT method that recycles trained adapters by finding a shared principal subspace. This allows for more efficient training of new adapters and inference with multiple adapters. Through experiments, we showed that EigenLoRAx works and provides practical benefits in a wide range of scenarios. Our method has the potential to mitigate the perpetually widening compute resource gap (Ahmed & Wahed, 2020; Besiroglu et al., 2024) and reduce the environmental cost of training and using AI models (Wu et al., 2021; Ligozat et al., 2021). It also holds promise for training personalized models (Tan et al., 2024) on low-resource devices, in privacy-critical use-cases. We have a large number of experiments ($6+$) on a diverse set of complex models, tasks and modalities. We have shown that EigenLoRAx excels in faster and efficient learning, memory savings and zero shot performance which differentiates it from conventional PeFT models. While our work is application-focused, we believe it is the first work to hypothesize and empirically prove the existence of shared weight subspaces of neural networks. This important insight has significant implications of model merging, efficiency, mechanistic interpretability, and neural learning theory.
@@ A.2 Natural Language Processing - GLUE benchmark
Table 11: Hyperparameters for EigenLoRAx for the GLUE benchmark. (Wang et al., 2019).
(RLrP - ReduceLRonPlateau)

### v3

@@ 3.2 Algorithm
Theorem 3.6 provides an upper bound on the Frobenius norm of the difference between $W_{d+1}$ or $W_{\mathcal{E}}$ and the optimal solution ${W^{*}}_{d+1}$. 2 provides a tighter upper bound on the norm of the difference when task $t_{d+1}$ majorly lies in the shared principal subspace. The extent to which task $t_{d+1}$ lies in the shared principle subspace is captured by the second term involving the sum of squared truncated singular values $\hat{W}$. Hence, if the task completely or majorly lie in the shared principal subspace, then the first term (sqrt(rank)) will dominate the upper bound. Hence, if rank($W_{d+1}\geq K$), then we can see that the upper bound in eq. 2 will be tighter than in eq. 1 where the task lies majorly in the shared principal subspace. Similarly, when $m\leq K$, the upper bound on the difference norm will be tighter for $W_{d+1}$ than $W_{\mathcal{E}}$. When ${W^{*}}_{d+1}$ has a significant alignment or projection along the singular vectors orthogonal to the ones with top $K$ singular values, then the second term in 1 comes into picture and it becomes difficult to directly compare the bounds in 1 and 2. However, if majority of the variance of ${W^{*}}_{d+1}$ is along the singular vectors orthogonal to the top $K$ components, it follows that $W_{\mathcal{E}}$ will never be able to achieve convergence while $W_{d+1}$. In contrast, $W_{d+1}$ could perform significantly better, as it is not restricted to learning only along the top $K$ principal components of $\hat{W}$. While the assumption that ${W^{*}}_{d+1}$ is spanned by the principal components of the shared principal subspace might appear to be very strong, we empirically observe in table 4 that such an assumption is not impractically far from reality. Particularly, we observe in table 2 that for GLUE benchmark, LoRA adapters trained on 5 diverse tasks shared a principal subspace. We see that EigenLoRAx was able to leverage the principal components of this shared subspace with just 12K training parameters learned for a new 6th task and achieve competitive performance compared to fine-tuning full rank weights with 125M parameters or individual LoRA adaptors with 1.2M parameters, even outperforming them in certain tasks. Similarly, table 4 demonstrates zeroshot performance using only top $K$ principal components of the shared subspace obtained through 500 LoRA adaptors trained on diverse tasks. This further suggests that increasing the number of LoRA adapters enables a richer set of top principal components, effectively spanning the shared subspace and providing broader coverage for new tasks.
@@
Memory Efficiency and Complexity
Our method demonstrates significant memory efficiency across experiments. A single set of EigenLoRAx PCs, combined with lightweight task-specific coefficients, can effectively replace both past and future LoRAs within a task domain. This is particularly advantageous when serving a large number of adapters, where frequent loading and unloading in VRAM incurs high latency, or keeping all adapters in memory demands excessive VRAM. For $d$ LoRAs of rank $r$ and $l$ layers, the memory footprint is $O(2drln)$. For EigenLoRAxs, it is $O(2Kl(d+n))$. As $r,K\ll n$, EigenLoRAx becomes more memory efficient in terms of memory required to save the models as $d$ increases. This becomes significantly useful for edge devices and large scale user serving AI systems.
4 Experiments and Results
@@ 5 Conclusion
We introduce EigenLoRAx, a significantly efficient model finetuning and inference method that recycles publicly available pretrained adapters by finding a shared principal subspace. This allows finetuning on new data by simply learning the lightweight coefficients of the shared subspace, and also requires less number of parameters to be saved for new tasks. Our comprehensive and diverse experiments show that EigenLoRAx is applicable to a large range of problems and model architectures. We believe that EigenLoRAx has the potential to mitigate the perpetually widening compute resource gap (Ahmed & Wahed, 2020; Besiroglu et al., 2024) and reduce the environmental cost of training and using machine learning models (Wu et al., 2021; Ligozat et al., 2021). It also holds promise for training personalized models (Tan et al., 2024) on low-resource devices, in privacy-critical use-cases. We have a large number of experiments ($6+$) on a diverse set of complex models, tasks and modalities. We have shown that EigenLoRAx excels in faster and efficient learning, memory savings and zero shot performance which differentiates it from conventional PeFT models. While our work is application-focused, we believe it is the first work to hypothesize and empirically prove the existence of shared weight subspaces of neural networks. This important insight has significant implications of model merging, efficiency, mechanistic interpretability, and neural learning theory.
A Note on Prior Work
This paper works on some ideas initially introduced in the EigenLoRA work (Kaushik et al., 2025), which was rejected from ICLR’25 and remains available on OpenReview. Due to irreconcilable differences regarding research ethics and authorship integrity, the primary authors of this work chose to independently develop and extend their contributions. These contributions in the previous draft included the initial algorithm, all experimental results (excluding Section 4.2.2), as well as all ablation studies and additional experiments. The code used for these experiments was also developed solely by the first authors of this work and is publicly available. Since the prior work was neither accepted nor published, there are no concerns regarding self-plagiarism. We also acknowledge discussions and exchanges with some of the earlier authors, which, while not directly contributing to the technical developments of this work, provided context and background that shaped our understanding of the broader problem. We have taken extensive measures to ensure that this paper exclusively reflects the contributions of the authors listed, and any similarities beyond our own contributions are purely coincidental.
@@ A.2 Natural Language Processing - GLUE benchmark
Table 11: Hyperparameters for EigenLoRAx for the GLUE benchmark. (Wang et al., 2019). (RLrP - ReduceLRonPlateau)